# ProbeWalk: Fast Estimation of Biharmonic Distance on Graphs via Probe-Driven Random Walks

### v1

@@ Abstract.
copyright: nonejournalyear: 2026copyright: noneccs: Networks Network algorithmsccs: Theory of computation Randomness, geometry and discrete structuresccs: Information systems Data mining
@@ 1. INTRODUCTION
In network analysis, graph distance metrics constitute foundational primitives, and numerous variants have been introduced to serve diverse tasks (yang2022Efficient; lü2011Link; shimada2016Graph; tsitsulin2020Just; jin2019Forest; tran2019Signed; zhang2019Pruning). An important example is the geodesic distance, which measures distance based on the shortest path between node pairs (newman2018Networks). Another notable metric is the resistance distance, arising from the theory of electrical networks, which has proved valuable in diverse domains including circuit theory, chemical graph applications, combinatorial matrix analysis, and spectral graph theory (doyle1984Random; klein1993Resistance; bapat2014Graphs; yang2013recursion; 2007Resistance).
@@
Among distance measures, the biharmonic distance has emerged as a versatile tool with demonstrated impact in clustering, learning, and network analysis (yi2018Biharmonica; yi2022Biharmonic; verma2017Hunt; kreuzer2021Rethinking; black2023Understandinga; tyloo2018Robustness; fan2020Spectrala; fitch2016Joint; xu2022Coherence). The biharmonic distance was originally proposed by Lipman et al. (lipman2010Biharmonic) for curved-surface geometry, a challenge in computer graphics and geometric processing. Prior work has studied the connection between the biharmonic distance and other measures. For example, yi2018Biharmonic connected the biharmonic distance to edge centrality and wei2021Biharmonic explored its relationship with the Kirchhoff index.
@@
The biharmonic distance balances neighborhood information with global spectral structure, providing a spectrum-sensitive alternative to geodesic and resistance distances across a range of tasks. Its versatility is exemplified in second-order noisy consensus, where performance metrics naturally scale with the squared distance (bamieh2012Coherence; yi2022Biharmonic). Beyond dynamical robustness, the biharmonic distance has found applications in other areas such as graph learning (kreuzer2021Rethinking; black2023Understandinga), physics (tyloo2018Robustness), and other network-related fields (fan2020Spectrala; fitch2016Joint; tyloo2019key).
@@
Despite its importance, computing the biharmonic distance remains a challenging task. Yi et al. (yi2018Biharmonic) proposed an algorithm for approximating biharmonic distances for all edges, but the algorithm needs to construct a large, dense matrix during preprocessing, which limits its applicability to real-world networks. Recently, Liu et al. (liu2024Fast) developed several efficient algorithms to estimate pairwise biharmonic distance based on random-walk sampling, and one of them, SWF, is the state-of-the-art algorithm. However, the limitation of this algorithm is that on large graphs, it requires a large number of trials to achieve good estimation accuracy, thus its runtime scales as $O(L^{5}/\varepsilon_{\mathrm{abs}}^{2})$. In practice, the truncation length $L$ tends to be large on real-world networks, and thus the $L^{5}$ term in the time complexity results in substantial computational overhead.
@@ 2.1. Notations
We work with the simple random walk on $\mathcal{G}$. Its transition matrix is $\bm{\mathit{P}}=\bm{\mathit{D}}^{-1}\bm{\mathit{A}}$, so $\bm{\mathit{P}}_{ij}=\frac{1}{\bm{\mathit{d}}_{i}}$ if $\{i,j\}\in E$ and $\bm{\mathit{P}}_{ij}=0$ otherwise. For $\ell\geq 0$, let $(\bm{\mathit{P}}^{\ell})_{ij}$ be the probability that a random walk from node $i$ visits node $j$ at the $\ell$-th hop. In this paper, following the convention, we assume $\mathcal{G}$ is not bipartite (which is the case for most real-world networks). According to (motwani1995Randomized), the random walks over $\mathcal{G}$ are ergodic, i.e., $\lim_{\ell\rightarrow\infty}(\bm{\mathit{P}}^{\ell})_{ij}=\bm{\pi}_{j}=\frac{\bm{\mathit{d}}_{j}}{2m}$ for any $i,j\in V$, where $\bm{\pi}$ denotes the stationary distribution of a random walk starting from any node. Writing $\Pi=\mathrm{diag}(\bm{\pi})$, the detailed balance holds: $\Pi\bm{\mathit{P}}=\bm{\mathit{P}}^{\top}\Pi$ (levin2017Markov). Let $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ be the eigenvalues of $\bm{\mathit{P}}$ with the column eigenvectors $\bm{\mathit{w}}_{1},\bm{\mathit{w}}_{2},\ldots,\bm{\mathit{w}}_{n}$, i.e., $\bm{\mathit{P}}\bm{\mathit{w}}_{j}=\lambda_{j}\bm{\mathit{w}}_{j}$. Let the mixing factor $\lambda=\max\{|\lambda_{2}|,|\lambda_{n}|\}$. Notably, $\lambda_{1}=1$ and $\bm{\mathit{w}}_{1}=\mathbf{1}$ (haveliwala2003Second).
@@ 2.2. Problem Definition
Since its introduction by Lipman et al. (lipman2010Biharmonic), the biharmonic distance on graphs has evolved into a broadly studied spectral distance (fitch2016Joint; yi2018Biharmonica; yi2018Biharmonic; yi2022Biharmonic; zhang2020Fast; black2023Understandinga; 2007Resistance; wei2021Biharmonic). In intuitive terms, a smaller biharmonic distance indicates two nodes are tightly connected through many routes, whereas a larger biharmonic distance reflects a more indirect connection. Its formal definition is as follows.
@@
According to (black2024Biharmonic), the biharmonic distance between any distinct nodes $s$ and $t$ satisfies $n^{-2}\leq\beta(s,t)\leq n^{3}$. This shows that the numerical values of the biharmonic distance can vary drastically across different node pairs. To provide a concrete illustration of this dynamic range, we perform a preliminary study on six real-world networks: for each network, we uniformly sample $100$ node pairs and compute their biharmonic distances by running the PUSH algorithm (liu2024Fast) for 1000 iterations (achieving a residual below $10^{-6}$). In Table 1, we show the minimum, maximum, and mean over 100 uniformly sampled node pairs per network.
@@ 2.3. Existing Algorithms
Building on Laplacian solvers and random projections, AppxBDRC approximates biharmonic distance for all edges of a graph $\mathcal{G}$ in time $O(m\log n/\varepsilon_{\mathrm{abs}}^{2})$ (yi2018Biharmonic). However, its preprocessing constructs a $(24\log n/\varepsilon_{\mathrm{abs}}^{2})\times n$ embedding matrix for a given $\varepsilon_{\mathrm{abs}}$, which itself requires $\tilde{O}(m/\varepsilon_{\mathrm{abs}}^{2})$ time and a full pass over the graph. This cost is undesirable when one only needs biharmonic distances for a small set of critical pairs. Moreover, the method is tailored to edge pairs; ad-hoc queries for arbitrary non-edge node pairs are not supported.
@@
liu2024Fast proposed an alternative expression of the biharmonic distance and its truncation bound.
Lemma 2.2 (Alternative Expression (liu2024Fast)).
@@
Lemma 2.3 (Truncation Bound (liu2024Fast)).
@@ 3.1. New Formula for Biharmonic Distance
We first introduce the Paired-Sign Probe. Inspired by Rademacher vectors (achlioptas2003Databasefriendly) and the antithetic-variates idea in Monte Carlo (robert2004Monte), the Paired-Sign Probe first zeros out $r$ entries to fix parity, then draws a uniformly random perfect matching on the remaining indices and assigns opposite signs within each pair to enforce pairwise cancellation. Unlike i.i.d. Rademacher probes, for which $\mathbb{E}[zz^{\top}]=I$, our probe matches a prescribed second moment $\mathbb{E}[zz^{\top}]=H$.
@@ 3.2. Random-Walk Based Estimator
We then construct a second-order U-statistic (hoeffding1948Class; serfling1980Approximation) from $R$ i.i.d. copies of $\widehat{\phi}(\bm{\mathit{z}})$ to obtain an unbiased estimator of $\phi(\bm{\mathit{z}})^{2}$. A straightforward idea is to approximate $\bm{\mathbb{E}}[\phi(\bm{\mathit{z}})^{2}]$ by averaging the squared probes from $R$ independent replicas of Eq. (9). However, this plug-in approach leads to bias: conditioning on $\bm{\mathit{z}}$, the identity
@@
To remove this variance while keeping the computation simple and streaming-friendly, we use the second-order U-statistic (hoeffding1948Class; serfling1980Approximation) built from $R$ i.i.d. replicas $\widehat{\phi}^{(1)}(\bm{\mathit{z}}),\ldots,\widehat{\phi}^{(R)}(\bm{\mathit{z}})$ of Eq. (9), all conditioned on the same $\bm{\mathit{z}}$:
@@
To obtain robustness without sacrificing simplicity, we use the median-of-means (MoM) estimator (nemirovsky1983Problem). According to (lugosi2019Mean), median-of-means delivers sub-Gaussian-style deviation bounds up to absolute constants and is robust to a minority of contaminated or unlucky blocks.
@@ 3.3. Algorithm Description
Fix a probe index $k$ and draw a seedable permutation $\pi_{k}:[n]\!\to\![n]$. We write $\pi_{k}^{-1}$ for its inverse permutation. Both $\pi_{k}$ and $\pi_{k}^{-1}$ can be evaluated in $O(1)$ time by instantiating them with a small-round Feistel pseudorandom permutation (luby1988How; black2002Ciphers).
@@ 23 $\bar{Y}_{g}\leftarrow\frac{1}{B}\sum_{k\in\mathcal{B}_{g}}Q_{R}(\bm{\mathit{z}}^{(k)})$
where $p$ is a fixed prime $\lfloor n/2\rfloor$. The parameters $a_{k}\in\{1,\dots,p-1\}$ and $b_{k}\in\{0,\dots,p-1\}$ are sampled from a seed deterministically derived from $k$ (e.g., via a keyed PRF (luby1988How; black2002Ciphers)).
@@
It has been proved in (li2023New) that
@@ 4.1. Experimental Setup
Datasets, query sets, and ground truth. Our experiments are conducted on nine real-world networks from SOMAR and SNAP (leskovec2014SNAP; hemphill2018Social), details of which are provided in Table 2. For each network, we uniformly sample 100 node pairs as the query set. On the first six networks, we compare ProbeWalk with existing algorithms to evaluate efficiency and accuracy; ground-truth biharmonic distances for the sampled pairs are computed by running PUSH for 1000 iterations (achieving a residual below $10^{-6}$(liu2024Fast). On the remaining three large-scale networks with tens of millions of nodes, we assess the scalability of our proposed algorithm; other algorithms are infeasible at this scale, so the error control is ensured by the theory of ProbeWalk together with our stabilization rule.
@@
Implementation details. All experiments are conducted on a Linux machine with an AMD EPYC 7713 (64-core, 2.0 GHz) processor and 1 TB of RAM. For fairness and reproducibility, we disable intra-query parallelism and execute each query on a single CPU thread with a uniform memory budget of 32 GB; no multi-threading or multi-processing is used within a query. Each experiment loads the graph into memory before timing begins. For the first six networks, the spectral quantities $\lambda_{2}$ and $\lambda_{n}$ are approximated via ARPACK (lehoucq1998ARPACK). All algorithms are implemented in Python. For randomized algorithms, we set the failure probability $\delta=0.01$, and we exclude a method if it fails to report the result for any query within 24 hours.
@@ 4.2. Query Efficiency on Real-World Networks
In this experiment, we aim to assess the efficiency and accuracy of ProbeWalk compared to PUSH, STW and SWF (liu2024Fast). We exclude AppxBDRC as a baseline because prior work has shown that PUSH, STW and SWF outperform it by several orders of magnitude.
@@ 4.3. Query Scalability on Large-Scale Networks
We fix the target relative error $\varepsilon=0.01$. The truncation length $L$ in Theorem 3.8 depends on the spectral quantity $\lambda=\max\{|\lambda_{2}|,|\lambda_{n}|\}$; estimating $\lambda$ on these graphs is infeasible within a one-day budget (lehoucq1998ARPACK). Instead, we select $L$ via the stabilization rule: starting at $L_{0}=1$, we double $L$ until the estimate stabilizes, and conservatively accept $L_{\mathrm{stop}}=2L$ if

### v3

@@ Abstract.
copyright: nonejournalyear: 2026copyright: noneconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesconference: ; -=$c$s: Networks Network algorithmsccs: Theory of computation Randomness, geometry and discrete structuresccs: Information systems Data mining
@@ 1. INTRODUCTION
In network analysis, graph distance metrics constitute foundational primitives, and numerous variants have been introduced to serve diverse tasks (Yang, 2022; lü2011Link; Shimada et al., 2016; Tsitsulin et al., 2020; Jin et al., 2019; Tran et al., 2019; Zhang et al., 2019). An important example is the geodesic distance, which measures distance based on the shortest path between node pairs (Newman, 2018). Another notable metric is the resistance distance, arising from the theory of electrical networks, which has proved valuable in diverse domains including circuit theory, chemical graph applications, combinatorial matrix analysis, and spectral graph theory (Doyle and Snell, 1984; Klein and Randić, 1993; Bapat, 2014; Yang and Klein, 2013; Chen and Zhang, 2007).
@@
Among distance measures, the biharmonic distance has emerged as a versatile tool with demonstrated impact in clustering, learning, and network analysis (Yi et al., 2018b, 2022; Verma and Zhang, 2017; Kreuzer et al., 2021; Black et al., 2023; Tyloo et al., 2018; Fan et al., 2020; Fitch and Leonard, 2016; Xu et al., 2022). The biharmonic distance was originally proposed by Lipman et al. (Lipman et al., 2010) for curved-surface geometry, a challenge in computer graphics and geometric processing. Prior work has studied the connection between the biharmonic distance and other measures. For example, Yi et al. (2018a) connected the biharmonic distance to edge centrality and Wei et al. (2021) explored its relationship with the Kirchhoff index.
@@
The biharmonic distance balances neighborhood information with global spectral structure, providing a spectrum-sensitive alternative to geodesic and resistance distances across a range of tasks. Its versatility is exemplified in second-order noisy consensus, where performance metrics naturally scale with the squared distance (Bamieh et al., 2012; Yi et al., 2022). Beyond dynamical robustness, the biharmonic distance has found applications in other areas such as graph learning (Kreuzer et al., 2021; Black et al., 2023), physics (Tyloo et al., 2018), and other network-related fields (Fan et al., 2020; Fitch and Leonard, 2016; Tyloo et al., 2019).
@@
Despite its importance, computing the biharmonic distance remains a challenging task. Yi et al. (Yi et al., 2018a) proposed an algorithm for approximating biharmonic distances for all edges, but the algorithm needs to construct a large, dense matrix during preprocessing, which limits its applicability to real-world networks. Recently, Liu et al. (Liu et al., 2024) developed several efficient algorithms to estimate pairwise biharmonic distance based on random-walk sampling, and one of them, SWF, is the state-of-the-art algorithm. However, the limitation of this algorithm is that on large graphs, it requires a large number of trials to achieve good estimation accuracy, thus its runtime scales as $O(L^{5}/\varepsilon_{\mathrm{abs}}^{2})$. In practice, the truncation length $L$ tends to be large on real-world networks, and thus the $L^{5}$ term in the time complexity results in substantial computational overhead.
@@ 2.1. Notations
We work with the simple random walk on $\mathcal{G}$. Its transition matrix is $\bm{\mathit{P}}=\bm{\mathit{D}}^{-1}\bm{\mathit{A}}$, so $\bm{\mathit{P}}_{ij}=\frac{1}{\bm{\mathit{d}}_{i}}$ if $\{i,j\}\in E$ and $\bm{\mathit{P}}_{ij}=0$ otherwise. For $\ell\geq 0$, let $(\bm{\mathit{P}}^{\ell})_{ij}$ be the probability that a random walk from node $i$ visits node $j$ at the $\ell$-th hop. In this paper, following the convention, we assume $\mathcal{G}$ is not bipartite (which is the case for most real-world networks). According to (Motwani and Raghavan, 1995), the random walks over $\mathcal{G}$ are ergodic, i.e., $\lim_{\ell\rightarrow\infty}(\bm{\mathit{P}}^{\ell})_{ij}=\bm{\pi}_{j}=\frac{\bm{\mathit{d}}_{j}}{2m}$ for any $i,j\in V$, where $\bm{\pi}$ denotes the stationary distribution of a random walk starting from any node. Writing $\Pi=\mathrm{diag}(\bm{\pi})$, the detailed balance holds: $\Pi\bm{\mathit{P}}=\bm{\mathit{P}}^{\top}\Pi$ (Levin et al., 2017). Let $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ be the eigenvalues of $\bm{\mathit{P}}$ with the column eigenvectors $\bm{\mathit{w}}_{1},\bm{\mathit{w}}_{2},\ldots,\bm{\mathit{w}}_{n}$, i.e., $\bm{\mathit{P}}\bm{\mathit{w}}_{j}=\lambda_{j}\bm{\mathit{w}}_{j}$. Let the mixing factor $\lambda=\max\{|\lambda_{2}|,|\lambda_{n}|\}$. Notably, $\lambda_{1}=1$ and $\bm{\mathit{w}}_{1}=\mathbf{1}$ (Haveliwala and Kamvar, 2003).
@@ 2.2. Problem Definition
Since its introduction by Lipman et al. (Lipman et al., 2010), the biharmonic distance on graphs has evolved into a broadly studied spectral distance (Fitch and Leonard, 2016; Yi et al., 2018b, a, 2022; Zhang et al., 2020; Black et al., 2023; Chen and Zhang, 2007; Wei et al., 2021). In intuitive terms, a smaller biharmonic distance indicates two nodes are tightly connected through many routes, whereas a larger biharmonic distance reflects a more indirect connection. Its formal definition is as follows.
@@
According to (Black et al., 2024), the biharmonic distance between any distinct nodes $s$ and $t$ satisfies $n^{-2}\leq\beta(s,t)\leq n^{3}$. This shows that the numerical values of the biharmonic distance can vary drastically across different node pairs. To provide a concrete illustration of this dynamic range, we perform a preliminary study on six real-world networks: for each network, we uniformly sample $100$ node pairs and compute their biharmonic distances by running the PUSH algorithm (Liu et al., 2024) for 1000 iterations (achieving a residual below $10^{-6}$). In Table 1, we show the minimum, maximum, and mean over 100 uniformly sampled node pairs per network.
@@ 2.3. Existing Algorithms
Building on Laplacian solvers and random projections, AppxBDRC approximates biharmonic distance for all edges of a graph $\mathcal{G}$ in time $O(m\log n/\varepsilon_{\mathrm{abs}}^{2})$ (Yi et al., 2018a). However, its preprocessing constructs a $(24\log n/\varepsilon_{\mathrm{abs}}^{2})\times n$ embedding matrix for a given $\varepsilon_{\mathrm{abs}}$, which itself requires $\tilde{O}(m/\varepsilon_{\mathrm{abs}}^{2})$ time and a full pass over the graph. This cost is undesirable when one only needs biharmonic distances for a small set of critical pairs. Moreover, the method is tailored to edge pairs; ad-hoc queries for arbitrary non-edge node pairs are not supported.
@@
Liu et al. (2024) proposed an alternative expression of the biharmonic distance and its truncation bound.
Lemma 2.2 (Alternative Expression (Liu et al., 2024)).
@@
Lemma 2.3 (Truncation Bound (Liu et al., 2024)).
@@ 3.1. New Formula for Biharmonic Distance
We first introduce the Paired-Sign Probe. Inspired by Rademacher vectors (Achlioptas, 2003) and the antithetic-variates idea in Monte Carlo (Robert and Casella, 2004), the Paired-Sign Probe first zeros out $r$ entries to fix parity, then draws a uniformly random perfect matching on the remaining indices and assigns opposite signs within each pair to enforce pairwise cancellation. Unlike i.i.d. Rademacher probes, for which $\mathbb{E}[zz^{\top}]=I$, our probe matches a prescribed second moment $\mathbb{E}[zz^{\top}]=H$.
@@ 3.2. Random-Walk Based Estimator
We then construct a second-order U-statistic (Hoeffding, 1948; Serfling, 1980) from $R$ i.i.d. copies of $\widehat{\phi}(\bm{\mathit{z}})$ to obtain an unbiased estimator of $\phi(\bm{\mathit{z}})^{2}$. A straightforward idea is to approximate $\bm{\mathbb{E}}[\phi(\bm{\mathit{z}})^{2}]$ by averaging the squared probes from $R$ independent replicas of Eq. (9). However, this plug-in approach leads to bias: conditioning on $\bm{\mathit{z}}$, the identity
@@
To remove this variance while keeping the computation simple and streaming-friendly, we use the second-order U-statistic (Hoeffding, 1948; Serfling, 1980) built from $R$ i.i.d. replicas $\widehat{\phi}^{(1)}(\bm{\mathit{z}}),\ldots,\widehat{\phi}^{(R)}(\bm{\mathit{z}})$ of Eq. (9), all conditioned on the same $\bm{\mathit{z}}$:
@@
To obtain robustness without sacrificing simplicity, we use the median-of-means (MoM) estimator (Nemirovsky and Yudin, 1983). According to (Lugosi and Mendelson, 2019), median-of-means delivers sub-Gaussian-style deviation bounds up to absolute constants and is robust to a minority of contaminated or unlucky blocks.
@@ 3.3. Algorithm Description
Fix a probe index $k$ and draw a seedable permutation $\pi_{k}:[n]\!\to\![n]$. We write $\pi_{k}^{-1}$ for its inverse permutation. Both $\pi_{k}$ and $\pi_{k}^{-1}$ can be evaluated in $O(1)$ time by instantiating them with a small-round Feistel pseudorandom permutation (Luby and Rackoff, 1988; Black and Rogaway, 2002).
@@ 23 $\bar{Y}_{g}\leftarrow\frac{1}{B}\sum_{k\in\mathcal{B}_{g}}Q_{R}(\bm{\mathit{z}}^{(k)})$
where $p$ is a fixed prime $\lfloor n/2\rfloor$. The parameters $a_{k}\in\{1,\dots,p-1\}$ and $b_{k}\in\{0,\dots,p-1\}$ are sampled from a seed deterministically derived from $k$ (e.g., via a keyed PRF (Luby and Rackoff, 1988; Black and Rogaway, 2002)).
@@
It has been proved in (Li and Sachdeva, 2023) that
@@ 4.1. Experimental Setup
Datasets, query sets, and ground truth. Our experiments are conducted on nine real-world networks from SOMAR and SNAP (Leskovec and Krevl, 2014; Hemphill et al., 2018), details of which are provided in Table 2. For each network, we uniformly sample 100 node pairs as the query set. On the first six networks, we compare ProbeWalk with existing algorithms to evaluate efficiency and accuracy; ground-truth biharmonic distances for the sampled pairs are computed by running PUSH for 1000 iterations (achieving a residual below $10^{-6}$(Liu et al., 2024). On the remaining three large-scale networks with tens of millions of nodes, we assess the scalability of our proposed algorithm; other algorithms are infeasible at this scale, so the error control is ensured by the theory of ProbeWalk together with our stabilization rule.
@@
Implementation details. All experiments are conducted on a Linux machine with an AMD EPYC 7713 (64-core, 2.0 GHz) processor and 1 TB of RAM. For fairness and reproducibility, we disable intra-query parallelism and execute each query on a single CPU thread with a uniform memory budget of 32 GB; no multi-threading or multi-processing is used within a query. Each experiment loads the graph into memory before timing begins. For the first six networks, the spectral quantities $\lambda_{2}$ and $\lambda_{n}$ are approximated via ARPACK (Lehoucq et al., 1998). All algorithms are implemented in Python. For randomized algorithms, we set the failure probability $\delta=0.01$, and we exclude a method if it fails to report the result for any query within 24 hours.
@@ 4.2. Query Efficiency on Real-World Networks
In this experiment, we aim to assess the efficiency and accuracy of ProbeWalk compared to PUSH, STW and SWF (Liu et al., 2024). We exclude AppxBDRC as a baseline because prior work has shown that PUSH, STW and SWF outperform it by several orders of magnitude.
@@ 4.3. Query Scalability on Large-Scale Networks
We fix the target relative error $\varepsilon=0.01$. The truncation length $L$ in Theorem 3.8 depends on the spectral quantity $\lambda=\max\{|\lambda_{2}|,|\lambda_{n}|\}$; estimating $\lambda$ on these graphs is infeasible within a one-day budget (Lehoucq et al., 1998). Instead, we select $L$ via the stabilization rule: starting at $L_{0}=1$, we double $L$ until the estimate stabilizes, and conservatively accept $L_{\mathrm{stop}}=2L$ if